OpenStudy (anonymous):

Drew is making a cardboard cut out of a right triangle. He knows that the longest side must be 15 inches long. One of the short sides must be 8 inches. How many inches long will he make the other side? Use the appropriate formula to find the solution and round to the nearest tenth. Show all steps or reasoning used. Be sure to check your answer for reasonableness.

OpenStudy (anonymous):

use pyth theorem. longest side is the hypothenuse. draw a pixx if you can not see and label, it does help. so if you would like to set it up I will see if yoou are on the right path

OpenStudy (anonymous):

I just wanted the formula mainly

OpenStudy (anonymous):

But could you stell me the steps without revealing the answer if possible

OpenStudy (anonymous):

*tell

OpenStudy (anonymous):

a^2 + b^2 = c^2 c^2 - b^2 = a^2 take the square root of both sides to get "a"

OpenStudy (anonymous):



OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

\[ x = \sqrt{15^2 - 8^2}\]

OpenStudy (anonymous):

thanks that helps a lot